# Towards Stable and Faithful Inpainting

### 08 Dec 2023 v1

@@ Abstract
Recent $*{}^{*}$: Equal contribution. ${}^{\dagger}$: Corresponding author. in inpainting increasingly relies on generative models, leveraging their strong generation capabilities for addressing ill-conditioned problems. However, this enhanced generation often introduces instability, leading to arbitrary object generation within masked regions. This paper proposes a balanced solution, emphasizing the importance of unmasked regions in guiding inpainting while preserving generative capacity. Our approach, Aligned Stable Inpainting with UnKnown Areas Prior (ASUKA), employs a reconstruction-based masked auto-encoder (MAE) as a stable prior. Aligned with the robust Stable Diffusion inpainting model (SD), ASUKA significantly improves inpainting stability. ASUKA further aligns masked and unmasked regions through an inpainting-specialized decoder, ensuring more faithful inpainting. To validate effectiveness across domains and masking scenarios, we evaluate on MISATO, a collection of several existing dataset. Results confirm ASUKA’s efficacy in both stability and fidelity compared to SD and other inpainting algorithms.
Figure 1: Comparison between ASUKA and reconstruction-based (MAE [20]) as well as generative-based (MAE-FAR [8], MAT [26], Co-Mod [61], LaMa [45], Stable Diffusion Inpainting Model [39]) inpainting algorithms. MAE offers stable masked region estimates, yet falls short in texture detail. GAN-based inpainting struggles with low fidelity, such as omitted white horizontal lines. SD is powerful but unstable, often introducing random elements and suffers from mask-unmask color inconsistency. ASUKA ensures consistency of masked-unmasked areas during diffusion and decoding processes, achieving stable and faithful inpainting. This image is from COCO [30].
@@ 1 Introduction
Image inpainting [5] generates content to fill in masked areas of images while maintaining consistency with the unmasked areas. Traditional inpainting algorithms [5, 14, 19, 25, 40] primarily rely on information from unmasked regions to reconstruct masked regions, resulting stable estimations but lacks fidelity. Moreover, the conventional reconstruction loss is prone to generating blurred synthesis due to the inherent ill-posed inpainting issues [36].
To facilitate inpainting with complex mask structures, recent approaches [35, 28, 7, 45, 46, 61] widely utilize the generative models [18, 22, 16], which introduce randomness into the inpainting process, like learning connections between image and noise distributions [18, 22] and utilizing Gibbs sampling [55, 11]. These induced stochasticity alleviates the ill-posed issues, yielding more flexible and faithful generation results [61, 62, 26].
However, every benefit comes with a cost. Current approaches mainly use random masking strategy to train inpainting algorithms, which usually leads to masking foreground objects. When trained with these masks, the model learns to generate foreground objects without information of the object, but sorely based on unrelated background. This introduces a bias in generative models, causing them to hallucinate unreasonable objects in the masked areas. This contradicts the goal of achieving stable inpainting.
In this paper, we propose to combine the strengths of both regression-based reconstruction and distribution-aligned generation to achieve stable and faithful inpainting simultaneously. Particularly, we adopt the Masked Auto-Encoder (MAE) [20] and Stable Diffusion inpainting model (SD) [39], respectively. MAE is trained simply to estimate masked pixels through the L2 reconstruction loss, resulting in stable and average results, albeit with some blurring. On the other hand, SD has impressive generation capacity, but may generate implausible content. In this paper, we demonstrate that the exceptionally faithful feature priors from MAEs can substantially enhance the inpainting stability of SD without sacrificing the performance as shown in Figure 1 with a frozen SD.
Particularly, we manipulate the generation and decoding processes of SD to achieve stable and faithful inpainting with the information from unmasked region.
i) In the generation process, we use MAE to estimate masked region sorely from unmasked region. Specifically, we remove the text condition and turn SD into a MAE prior guided image inpainting model. To align SD with MAE while keeping SD frozen, we propose an alignment module, which is trained on MAE-prior-to-image task with standard diffusion objective and achieves dimension and distribution alignment to cross-attention layers of SD.
ii) In the decoding process, we ensure consistency between the masked and unmasked regions. The use of Variational Auto-Encoder trained with KL divergence loss (KL-VAE) in SD inevitably leads to degradation, causing the color inconsistencies in some inpainting cases. Zhu et al. [64] observed similar issue and train a decoder conditioned on the unmasked regions. However, their decoder only preserves the unmasked region and still suffers from the inconsistency between the masked and unmasked regions. To tackle this challenge, we re-formulate the decoder of KL-VAE as a local harmonization model conditioned on unmasked image pixels. Such a specialized inpainting decoder significantly improves the perceptual consistency of inpainted images.
@@
Combined together, our proposed system, dubbed Aligned Stable inpainting with Unknown Areas prior (ASUKA), is illustrated in Figure 2. To evaluate the effectiveness of inpainting algorithms across various scenarios and mask shapes, we utilize an evaluation dataset named MISATO, which selects representative testing images from Matterport3D [10], Flickr-Landscape [29], MegaDept[27], and COCO 2014 [30]. This dataset thus covers four distinct domains—landscape, indoor, building, and background—making it diverse enough to serve as a benchmark for evaluation. Through experiments conducted on MISATO, especially with large irregular masks, we validate the efficacy of ASUKA.
Contributions. Our contributions can be summarized as:
  • We emphasize stability in inpainting tasks as the generative models suffers from the instability problem;
  • We balance fidelity and stability through aligning reconstruction-based Masked Auto-Encoder (MAE) and generation-based Stable Diffusion (SD).
  • We ensure mask-unmask consistency in the decoding process of SD by re-formulating decoding as a local harmonization task conditioned on unmasked image pixels and training a inpainting-specialized decoder.
@@ 2 Related Works
Image inpainting is a longstanding challenge to infill the missing regions with consistent pixels. Classical approaches, relying on patchmatch [15, 4, 58] and partial differential equation [5, 9, 6] are constrained by low-level image features and often fail to produce reasonable results for large holes. Benefits from the adversarial training [18], learning-based GANs achieve prominent success in image inpainting [36, 54, 61, 26, 8]. In particular, GAN-based inpainting methods have evolved with adaptive convolutions [31, 54, 57], attention mechanism [53, 52, 56, 24], and high-resolution generalization in frequency learning [45, 51, 13]. Methods like Co-Mod [61] prioritize adversarial learning and plausible generation to address the challenging ill-posed inpainting issue [26, 62], yielding robust and high-quality generations, even for large masks. However, they suffer from unstable results and uncontrollable hallucinations conditioned on different random latent codes. On the other hand, methods emphasizing higher reconstruction penalties [35, 45, 8] achieve more stable generations, but struggle with large masks, often resulting in blur artifacts. Recent diffusion models [41, 39] have rapidly developed and demonstrated impressive inpainting results. However, these models share similarities with GAN-based methods, learning to align distributions rather than achieving faithful pixel space reconstruction. As a consequence, these diffusion-based approaches face challenges in maintaining faithful inpainting results.
@@
Adapting Stable Diffusion. Currently, SD [39] stands out as one of the most widely used diffusion models. SD incorporates a KL-VAE to learn a latent space preserving image semantics at a lower resolution, along with a U-Net for diffusion within this latent space. Various tuning and aligning methods have been developed to introduce additional conditions to SD. This includes image-inversion for text-guided image-to-image translation [34], textual-inversion for personalization [17], LoRA fine-tuning [23], and controlnet [59] to add other types of conditions. In this paper, our objective is to enhance stability while preserving the generation capacity of the stable diffusion model. Consequently, we refrain from fine-tuning the U-Net backbone of SD. Conversely, the text-condition is not utilized in inpainting tasks. Textual-inversion [17] introduces learnable tokens in Clip text encoders to improve subject description. ObjectStitch [43] employs the CLIP visual feature to represent the subject. In our approach, we advocate removing the text-condition and instead guiding SD using the Masked Auto-Encoder (MAE) prior for masked region.
@@
Information loss in SD latent space. Although claimed only eliminates imperceptible details, the KL-VAE used by SD causes distortion in the reconstruction of images. OpenAI [2] proposes a larger decoder to improve the decoding quality of SD’s latent. Luo et al. [33] propose a frequency-augmented decoder to address the super-resolution case. Zhu et al. [64] propose to preserve unmasked regions during decoding. In this paper, inspired by these approaches, we take a step further to ensure the low-frequency semantic consistency between mask and unmask regions.
@@
Masked Image-Modeling [3] (MIM) is an active research area in self-supervised learning. Typical MIM methods [3, 20, 50, 12] split images into visible and masked patches, learning to estimate masked patches from visible patches. Training targets for visible patches encompass pixel values [20], HOG features [48], and high-level semantic features [49]. While the primary objective of MIM is representation learning, its potential effectiveness in image generation is also noteworthy. Cao et al. [8] adopts MAE features and attention scores to assist the convolutional inpainting model better in handling long-distance dependencies. In contrast, this paper proposes facilitating MAE as the inpainting prior for SD and enhancing the stability of the inpainting process by incorporating MAE prior guidance into the cross-attention modules of SD.
@@ 3 Methodology
Problem setup. Inpainting takes as inputs a masked image to complete with a mask to indicate the missing region. The target of inpainting is to fill the missing region based on the information of unmasked regions to generate high-fidelity images. In this paper, we focus on the standard inpainting task without utilizing other conditions.
Figure 2: ASUKA adopts the MAE to provide a low-resolution stable prior for frozen SD to maintain the generation capacity while increasing stability. A better decoder is utilized to achieve mask-unmask semantic consistency when decoding SD results from latent to image for better fidelity.
Overview. The framework of the proposed Aligned Stable inpainting with Unknown Areas prior (ASUKA) is illustrated in Figure 2. ASUKA adopts the pre-trained stable diffusion inpainting model (SD) [39]. Our target is to provide more stable and faithful inpainting results while fully exploiting the generation capacity of SD. Our alignment includes i) a stable alignment to align stable Masked Auto-Encoder (MAE) prior for masked region with SD and ii) a faithful alignment to align ground-truth unmasked region with generated masked region during decoding. To this end, we freeze the U-Net of SD, while replacing the text-condition part with our proposed MAE prior to increase stability. To align the MAE prior to SD, we introduce a alignment module, which is trained via the training objective of SD. Additionally, to align masked and unmasked regions during decoding and resolve the problem of information loss from the KL-VAE decoder which causes mask-unmask inconsistency to SD, we train an inpainting-specialized decoder to better decode the latent back to the image space for seamless fidelity. Combined together, ASUKA achieves more stable and faithful inpainting results.
@@ 3.1 Masked Auto-Encoder Prior
Figure 3: Use MAE prior for image-to-image translation (start from 80% noise rate) via SD achieves poor inpainting results.
@@
MAE as stable prior. While recent generative models rely on stochastic noise to provide more diverse generation results, it leads to the generation of random objects unexpectedly. On the other hand, MAE is known to provide a stable estimation of masked regions based purely on the unmasked regions. As MAE is trained on the L2 reconstruction loss, we can regard the estimation of MAE as a mean estimation, which can be utilized to provide a stable prior for generative models to not generate new concepts. However, MAE itself results in average and blur generations and cannot reconstruct detailed textures of the masked region, and works poorly if we use MAE prior as the initial values for the stable diffusion model to perform image-to-image style inpainting, as illustrated in Figure 3. To this end, we adopt the MAE to provide prior to stabilizing SD.
@@
Masking strategy for MAE. However, the original MAE is trained to estimate random masks uniformly distributed in the image. It is infeasible to directly adopt pre-trained MAE for inpainting task, which usually contains large continuous holes in the masked region. Inspired by Cao et al. [8], we fine-tune the MAE to inpainting masks. To adapt MAE for more practical inpainting scenarios, unlike Cao et al. [8] which only use continuous masks of range $[0.1,0.5]$, we construct a systematic masking strategy. Specifically, we generate masks with 50% object-shape mask, 20% Co-Mod mask [61], 20% LaMa [45] mask, 5% rectangle mask and 5% complement rectangle mask. We combine object-shape mask with Co-Mod style and LaMa style with the chance of 25%, respectively. This construction of mask style estimates the masks occurs in inpainting tasks, especially for the object removal and user-specified irregular masks. We control the mask ratio in the range of $[0.1,0.75]$ to follow the training scenario of MAE. For masks smaller than the ratio of 75%, we enlarge the mask ratio to 75% with randomly selected mask regions.
In practice, we train the MAE on images of size $256^{2}$, which is efficient and produce stable guidance for stable diffusion to generate high-resolution images.
3.2 Align MAE Prior with Stable Diffusion

### 17 Mar 2024 v2

# Project Page: https://yikai-wang.github.io/asuka/ MISATO Dataset: https://github.com/Yikai-Wang/asuka-misato

@@ Abstract
Recent progress in inpainting increasingly relies on generative models, leveraging their strong generation capabilities for addressing large irregular masks. However, this enhanced generation often introduces context-instability, leading to arbitrary object generation within masked regions. This paper proposes a balanced solution, emphasizing the importance of unmasked regions in guiding inpainting while preserving generation capacity. Our approach, Aligned Stable Inpainting with UnKnown Areas Prior (ASUKA), employs a Masked Auto-Encoder (MAE) to produce reconstruction-based prior. Aligned with the powerful Stable Diffusion inpainting model (SD), ASUKA significantly improves context stability. ASUKA further adopts an inpainting-specialized decoder, highly reducing the color inconsistency issue of SD and thus ensuring more visual-consistent inpainting. We validate effectiveness of inpainting algorithms on benchmark dataset Places 2 and a collection of several existing datasets, dubbed MISATO, across diverse domains and masking scenarios. Results on these benchmark datasets confirm ASUKA’s efficacy in both context-stability and visual-consistency compared to SD and other inpainting algorithms.
Keywords:
Image Inpainting Context-Stability Visual-Consistency
Figure 1: Comparison on $1024^{2}$ image (depicted from MegaDepth [26]) between our proposed ASUKA and reconstruction-based (MAE [19]) as well as generative-based (MAE-FAR [7], MAT [25], Co-Mod [59], LaMa [45], Stable Diffusion [39]) inpainting models. MAE offers stable estimates, yet falls short in texture detail. GAN-based inpainting struggles with low fidelity. Powerful SD hallucinates random elements and suffers from mask-unmask visual inconsistency (color of clouds in the orange square). ASUKA achieves context-stable and visual-consistent inpainting.
@@ 1 Introduction
Figure 2: Object removal on $1024^{2}$ images. ASUKA enjoys context-stable (first to second rows) and visual-consistent (second to third rows) inpainting compared with SD.
Image inpainting [5] fills masked areas of images while maintaining consistency with the unmasked regions. Traditional inpainting algorithms [5, 13, 18, 24, 40] often result in blurred synthesis when reconstructing masked regions, as indicated by [36]. The advent of Generative Adversarial Networks (GANs) has led to works [35, 27, 8, 45, 46, 59, 17, 21, 15] that utilize GANs to fill complex mask structures, achieving impressive inpainting results. However, they still struggle with general challenging inpainting cases, particularly in filling large holes.
The recent Stable Diffusion Inpainting (SD) [39] model, with its large model capacity and extensive dataset training, acts as a versatile tool for image inpainting. However, we have observed a phenomenon of "context-instability" problem, where the model generates random, unreasonable elements to fill masked regions of images, as depicted in first to second rows in Fig. 2, as well as Fig. 1 and Fig. 8. This issue appears to come from the random masking strategy used to train SD. This strategy introduces training cases where foreground objects are completely masked and the SD is forced to fill masked regions with foreground objects. Consequently, SD will hallucinate unreasonable objects devoid of contextual information. Adjusting prompts may reduce this risk, but the best prompt is image-dependent, making it infeasible for practical applications.
Furthermore, the inpainted results of SD may suffer from "visual inconsistency" problem. This problem, less explored in academia but critical for real-world industrial applications, results in significant color discrepancies between inpainted and unmasked regions and exhibits smear-like traces in the image, as shown in the second to third rows in Fig. 2, as well as Fig. 1 and Fig. 8. Essentially, this visual-inconsistency comes from the misalignment between the pixel distributions of filled results and original images due to information loss in the Kullback-Leibler Variational Autoencoder (KL-VAE) used in SD, as illustrated in Fig. 5. To improve visual consistency, an inpainting-specialized VAE decoder is necessary to explicitly mitigate information loss based on available ground-truth information of unmasked images.
To efficiently tackle context-instability and visual inconsistency issues, we enhance the SD [39] with regression-based reconstruction and distribution-aligned generation. This results in an improved image inpainting model that avoids generating unreasonable elements in the masked region and reduces mask-unmask color inconsistency. The vanilla SD [39] adopts a KL-VAE to compress image into latent and perform diffusion in the latent space. We manipulate its generation and decoding processes for better context stability and visual consistency. i) We propose using the Masked Auto-Encoder (MAE)[19] as a prior to guide and stabilize the SD’s generation process. As shown in Fig. 2, MAE yields stable yet blurred results, while SD may produce implausible content despite its impressive generation capacity. By aligning MAE prior with SD objectives, we improve context-stability without compromising performance. ii) We redesign the KL-VAE decoder in SD to address color inconsistencies between masked and unmasked regions. The new decoder, acting as a local harmonization model conditioned on unmasked image pixels, significantly enhances visual consistency.
Formally, to enhance the context-stability and visual-consistency of image inpainting model, we present the Aligned Stable Inpainting with Unknown Areas prior (ASUKA) framework. i) To enhance context-stability, we replace the text-condition part of the SD with the MAE prior. We fine-tune the MAE to ensure adaptability to practical inpainting scenarios, as MAE is trained on uniformly distributed masks but inpainting usually contains large continuous masks. We train an alignment module to align MAE prior with SD using the standard SD objective while freezing the fine-tuned MAE and SD. ii) To address visual-consistency issues during decoding, we redesign the decoder of the KL-VAE in SD. We inject additional inputs of unmasked image pixels and 0-1 masks to the decoder. This new decoder is trained to perform local harmonization task, taking as inputs of inharmonious latent and taking as outputs the original harmonious image. We generate inharmonious latent via the proposed color and latent augmentation techniques to incorporate KL-VAE information loss and generation artifacts of SD. Then the decoder is trained to recover the original image from corrupted latent based on the ground-truth information of unmasked image. These steps collectively enable ASUKA to achieve more context-stable and visual-consistent inpainting results.
@@
To evaluate the effectiveness of inpainting algorithms across various scenarios and mask shapes, in addition to the benchmark dataset Places 2 [61], we further utilize an evaluation dataset named MISATO, which selects representative testing images from Matterport3D [10], Flickr-Landscape [28], MegaDept[26], and COCO 2014 [29]. This dataset thus covers four distinct domains—landscape, indoor, building, and background—making it diverse enough to serve as a benchmark for evaluation. Experiments on MISATO and Places 2, especially with large irregular masks, validate the efficacy of ASUKA.
Contributions. ASUKA enhances image inpainting from context-stability and visual-consistency while leveraging the generation capacity of the frozen SD inpainting model. It achieves this through two main components: i) Context-Stable Alignment: ASUKA aligns the stable MAE prior with SD to provide a context-stable estimation of masked regions, replacing the text-condition part of SD with MAE prior. ii) Visual-Consistent Alignment: ASUKA re-formulates the decoding from SD latent to image as a local harmonization task, trains an inpainting-specialized decoder to align masked and unmasked regions during decoding and thus mitigates color inconsistencies caused by information loss from the SD.
@@ 2 Related Works
Image inpainting is a longstanding challenge to infill the missing regions with consistent pixels. Classical approaches, relying on patchmatch [14, 4, 56] and partial differential equation [5, 9, 6] are constrained by low-level image features and often fail to produce reasonable results for large holes. GAN-based [17] inpainting methods [36, 53, 59, 25, 7] have evolved with adaptive convolutions [30, 53, 54], attention mechanism [52, 51, 55, 23], and high-resolution generalization in frequency learning [45, 50, 12]. Methods like Co-Mod [59] prioritize adversarial learning and plausible generation to address the challenging ill-posed inpainting issue [25, 60], yielding robust and high-quality generations, even for large masks. However, they suffer from unstable results and uncontrollable hallucinations conditioned on different random latent codes. On the other hand, methods emphasizing higher reconstruction penalties [35, 45, 7] achieve more stable generations, but struggle with large masks, often resulting in blur artifacts. Recent diffusion models [41, 39] have rapidly developed and demonstrated impressive inpainting results. However, these models share similarities with GAN-based methods, learning to align distributions rather than achieving faithful pixel space reconstruction. As a consequence, diffusion-based approaches also suffer from context-instability issue.
@@
Adapting Stable Diffusion. SD [39] is one of the most widely used diffusion models. SD incorporates a KL-VAE to learn a latent space preserving image semantics at a lower resolution, along with a U-Net for diffusion within this latent space. Various tuning and aligning methods have been developed to introduce additional conditions to SD, including image-inversion for text-guided image-to-image translation [34], textual-inversion for personalization [16], LoRA fine-tuning [22], and controlnet [57] to add other types of conditions. Our objective is to enhance context-stability while preserving the generation capacity of SD. Consequently, we refrain from fine-tuning the U-Net backbone of SD. The text-condition is not utilized in inpainting tasks. Textual-inversion [16] introduces learnable tokens in Clip text encoders to improve subject description. ObjectStitch [43] employs the CLIP visual feature to represent the subject. In our approach, we advocate removing the text-condition and instead guiding SD using the Masked Auto-Encoder (MAE) prior for masked region.
@@
Information loss in SD latent space. Although claimed only eliminates imperceptible details, the KL-VAE used by SD causes distortion in the reconstruction of images. See Fig. 5(a) for illustrative examples. OpenAI [2] proposes a larger decoder to improve the decoding quality of SD’s latent. Luo et al. [33] propose a frequency-augmented decoder to address the super-resolution case. Zhu et al. [62] propose to preserve unmasked regions during decoding. In this paper, we ensure the low-frequency color visual-consistency in the decoding process of SD.
@@
Masked Image-Modeling [3] (MIM) is an active research area in self-supervised learning. Typical MIM methods [3, 19, 49, 11] split images into visible and masked patches, learning to estimate masked patches from visible patches. Training targets for visible patches encompass pixel values [19], HOG features [47], and high-level semantic features [48]. While the primary objective of MIM is representation learning, its potential effectiveness in image generation is also noteworthy. Cao et al. [7] adopts MAE features and attention scores to assist the convolutional inpainting model better in handling long-distance dependencies. In contrast, this paper uses MAE prior to enhance the context-stability of SD.
@@ 3 Methodology
Problem setup. Inpainting takes as inputs a masked image to complete with a mask to indicate the missing region. The target of inpainting is to fill the missing region based on the information of unmasked regions to generate high-fidelity images. In this paper, we focus on the standard inpainting task without utilizing other conditions. We adopt stable diffusion inpainting model [39] and focus on the general issues of inpainting models, i) context-instability issue: unstable and uncontrollable hallucinations, yielding random elements generated in the masked region; ii) visual-inconsistency issue: mask-unmask color inconsistency issue, yielding smear-like traces in the masked region.
Figure 3: ASUKA tackles the context-instability issue by adopting the MAE to provide a stable prior for frozen SD to maintain the generation capacity while increasing context-stability. For the visual-inconsistency issue, ASUKA utilizes an inpainting-specialized decoder to achieve mask-unmask color consistency when decoding SD latent.
(a) Use MAE prior for image-to-image translation (start from 80% noise rate) via SD achieves poor inpainting results.
(b) Alignment module.
Figure 4: (a) Use MAE as initial value v.s. as condition. (b) Alignment module.
Overview. The framework of the proposed Aligned Stable inpainting with Unknown Areas prior (ASUKA) is illustrated in Fig. 3. ASUKA adopts the pre-trained stable diffusion inpainting model (SD) [39]. Our target is to provide more context-stable and visual-consistent inpainting results while fully exploiting the generation capacity of SD. ASUKA includes i) a context-stable alignment to align stable Masked Auto-Encoder (MAE) prior for masked region with SD and ii) a visual-consistent alignment to align ground-truth unmasked region with generated masked region during decoding. To this end, we freeze the U-Net of SD, while replacing the text-condition part with our proposed MAE prior to increase context-stability. To align the MAE prior to SD, we introduce an alignment module, which is trained via the training objective of SD. Additionally, to align masked and unmasked regions during decoding and resolve the problem of information loss from the KL-VAE decoder which causes mask-unmask color inconsistency, we train an inpainting-specialized decoder to decode the latent back to the image space for seamless visual-consistency. Combined together, ASUKA achieves more context-stable and visual-consistent inpainting.
@@ 3.1 Masked Auto-Encoder Prior
MAE as context-stable prior. While recent generative models rely on random noise to provide more diverse generation results, it leads to the generation of random objects unexpectedly. In contrast, MAE is known to provide a context-stable estimation of masked regions based purely on the unmasked regions. As MAE is trained on the L2 reconstruction loss, we can regard the estimation of MAE as a mean estimation, which can be utilized to provide a context-stable prior for generative models to not generate new concepts. However, MAE itself results in average and blur generations and cannot reconstruct detailed textures of the masked region, and works poorly if we use MAE prior as the initial values for the SD to perform inpainting in image-to-image style, as illustrated in Fig. 3(a). To this end, we adopt the MAE to provide prior to stabilizing SD.
@@
Training MAE. The original MAE is trained to estimate random masks uniformly distributed in the image, while inpainting task usually contains large continuous masks. Inspired by Cao et al. [7], we fine-tune the MAE to inpainting masks. To adapt MAE for more practical inpainting scenarios, we construct a systematic masking strategy. The mask basis contains: object-shape mask, irregular mask, and regular mask. We collect object-shape masks from COCO [29] object segments. We use irregular masks from previous studies, including Co-Mod mask [59] and LaMa [45] mask. The regular masks contains rectangle and complement rectangle mask. To ensure generalization and coverage, for each mask we generate from mask basis with the probability of 50% object-shape, 40% irregular, and 10% regular. For object-shape mask basis, we combine it with irregular mask with the chance of 50%. This construction of mask style estimates the masks occurs in inpainting tasks, especially for the object removal and user-specified irregular masks. We control the mask ratio in the range of $[0.1,0.75]$ to follow the training scenario of MAE. For masks smaller than the ratio of 75%, we enlarge the mask ratio to 75% with randomly selected mask regions. This benefits ASUKA to tackle the large hole inpainting task.
3.2 Align MAE Prior with SD for Context-Stability